OpenStudy (anonymous):

Where does the graph of y = 2x2 − 5x + 3 cross the x-axis?

myininaya (myininaya):

when y is 0

OpenStudy (anonymous):

thts the answer?

jhonyy9 (jhonyy9):

x_1,_2=(5+/- sqrt(25-24))/4 =(5+/- 1)/4 = 4/4 =1 and 6/4 =3/2 - so will cross the x-axis in point 1 and 3/2